# Thompson Sampling for Repeated Newsvendor

### v1

@@
Abstract.
In this paper, we investigate the performance of Thompson Sampling (TS) for online learning with censored feedback, focusing primarily on the classic repeated newsvendor model–a foundational framework in inventory management–and demonstrating how our techniques can be naturally extended to a broader class of problems. We model demand using a Weibull distribution and initialize TS with a Gamma prior to dynamically adjust order quantities. Our analysis establishes optimal (up to logarithmic factors) frequentist regret bounds for TS without imposing restrictive prior assumptions. More importantly, it yields novel and highly interpretable insights on how TS addresses the exploration-exploitation trade-off in the repeated newsvendor setting. Specifically, our results show that when past order quantities are sufficiently large to overcome censoring, TS accurately estimates the unknown demand parameters, leading to near-optimal ordering decisions. Conversely, when past orders are relatively small, TS automatically increases future order quantities to gather additional demand information. Extensive numerical simulations further demonstrate that TS outperforms more conservative and widely-used approaches such as online convex optimization, upper confidence bounds, and myopic Bayesian dynamic programming. This study also lays the foundation for exploring general online learning problems with censored feedback.
conference: ; ;
1. Introduction
@@
More broadly, the repeated newsvendor problem serves as a key representative of a broader class of problems referred to as “online learning with censored feedback.” In this setting, the observation is always the minimum of an unknown random variable and the chosen action. For instance, in the newsvendor problem, the censored feedback corresponds to the minimum of the demand and the stock order. Similarly, in an auction, it is given by the minimum of the buyer’s willingness to pay and the seller’s set price. These problems inherently exhibit a trade-off between “large exploration”—choosing a sufficiently large action to better observe demand or willingness to pay for more accurate estimation—and “optimal exploitation”—making the right decision to minimize regret. While this paper primarily focuses on the repeated newsvendor problem, we also take an initial step toward systematically exploring the broader class of online learning problems with censored feedback.
Existing studies on the repeated newsvendor problem have established a $\sqrt{T}$-regret bound under fairly general unknown demand distributions and censored feedback, often leveraging the online convex optimization (OCO) framework (Huh and Rusmevichientong, 2009). However, as widely recognized in the bandit and online learning literature (Chapelle and Li, 2011; Seldin and Slivkins, 2014; Xu and Zeevi, 2023), TS often outperforms OCO-based approaches (which were originally developed for adversarial online learning) as well as other methods such as the Upper Confidence Bound (UCB). These advantages are supported by extensive numerical experiments and theoretical analyses in the aforementioned studies, as well as in our own work. This motivates us to adopt TS as the preferred approach for the repeated newsvendor problem and beyond.
1.1. Main Contributions
In this study, we provide a systematic analysis by establishing a $\sqrt{T}$-frequentist regret bound for TS in the repeated newsvendor problem. Our contributions are four-fold.
First, this study presents the first regret analysis of TS, one of the best-performing and most widely adopted algorithms in the bandit literature, in the repeated newsvendor model with censored feedback. This sets our work apart from prior studies that primarily focus on more conservative approaches such as OCO and UCB. Additionally, our frequentist regret analysis provides guarantees that hold for arbitrary underlying Weibull demand parameters or priors, further distinguishing it from Bayesian dynamic programming approaches that rely on restrictive assumptions, such as a well-specified Gamma prior. See Section 1.3 for details.
Second, our study provides a highly interpretable framework for analyzing the exploration-exploitation trade-off. Specifically, it offers insights into estimation error of unknow demand and enables a closed-form understanding of how TS naturally balances large-scale exploration with optimal exploitation in an intuitive and automatic manner. This novel perspective, to the best of our knowledge, provides the most interpretable explanation of the exploration-exploitation trade-off in the repeated newsvendor problem. A brief explanation is provided in the “Main Messages” subsection (Section 1.2), followed by a more detailed theoretical analysis in Section 3.
Third, through extensive numerical experiments, we demonstrate that TS outperforms existing approaches, as shown in Section 4. This finding aligns with the widely recognized effectiveness of TS in prior studies.
Fourth, TS and our analytical framework naturally extend to the broader settings of online learning with censored feedback, making it a versatile and effective approach for a wide range of decision-making problems under uncertainty. This extension requires only that the regret is Lipschitz continuous with respect to actions (Assumption 1) and that the relationship between the optimal action and the underlying parameter is continuous and monotone (Assumption 2). Further details and discussion of these assumptions are provided in Section 5.
1.2. Main Messages
In this paper, we investigate an online learning problem with censored feedback using the classic newsvendor model–one of the most fundamental frameworks in inventory management–as our pivotal example. Specifically, we consider a setting where the true demand $D_{t}$ is unknown, the action $y_{t}$ is the order quantity, and the observation $Y_{t}$ is censored feedback given by
(1) $Y_{t}=\min\{D_{t},y_{t}\}.$
@@ 1.2. Main Messages
Where demand is exactly observed when sales are less than the order quantity, that is, when $D_{t}<y_{t}$; and the demand is censored at the order quantity when sales equal $y_{t}$, that is, when $D_{t}\geq y_{t}$. The newsvendor setting experiences censored feedback–the decision-maker never observes lost-sales if demand exceeds the order quantity. This makes it difficult to accurately estimate demand, as it requires finding the right balance between not ordering too much to prevent excess inventory and placing larger orders to better understand how much demand is actually being missed (i.e., the afore-mentioned exploration-exploitation trade-off).
To address this trade-off, we model the demand distribution by a Weibull distribution–a flexible and widely used parametric family–and propose using TS to dynamically select order quantities. Our key insights include:
@@
Regardless of the algorithm used, we derive the confidence interval for demand estimation under censored feedback (1). The estimation error at round $t$ scales inversely with $\sum_{i=1}^{t-1}(1-e^{\theta^{*}y_{i}^{k}})$, where $\theta^{*}$ and $k$ are the scale and shape parameters of the Weibull distribution. This provides a rigorous quantification of how smaller past actions lead to larger errors and highlights the critical trade-off between large exploration and optimal exploitation.
@@
Empirical Effectiveness
@@
We conduct extensive numerical experiments demonstrating that TS yields competitive performance in terms of cumulative regret, outperforming existing widely-used approaches such as OCO, UCB, and myopic Bayesian dynamic programming. These experiments confirm the widely recognized belief on the effectiveness of TS in online learning and bandit literature and practical applications (Chapelle and Li, 2011; Seldin and Slivkins, 2014; Xu and Zeevi, 2023).
Extensions to Online Learning with Censored Feedback
We illustrate how our analytical framework naturally extends to broader settings of online learning with censored feedback, making it applicable to a wide range of problems where the feedback is censored. As validated by Assumptions 1 and 2 in Section 5, this extension requires only that the regret is Lipschitz continuous with respect to actions and that the relationship between the optimal action and the underlying parameter is continuous and monotone. We also discuss the technical limitation and possible refinement of these assumptions.
1.3. Related Work
This paper investigates online statistical learning and optimization in inventory control, specifically in a finite-horizon repeated newsvendor problem where the demand distribution parameters are initially unknown and must be learned over time. We focus on a perishable product with unobserved lost-sales, where sales data are censored by inventory levels, and any excess inventory does not carry over to the next period. The decision-maker must determine the order quantity before observing demand realization in that period. To address this challenge, we apply TS, a Bayesian approach that iteratively updates demand beliefs based on censored observations. This framework effectively balances exploration and exploitation, leading to improved inventory decisions over time and reducing long-term regret associated with demand uncertainty.
1.3.1. Bayeisan Dynamic Programming Literature
The first stream of research has formulated this problem using an offline dynamic programming (DP) approach, typically solved via backward induction. However, backward induction often suffers from the curse of dimensionality, making it computationally intractable for large-scale problems. Consequently, much of the existing literature in this area has focused on heuristic solutions as approximations to the optimal policy. (Chen, 2010) propose heuristics based on the bounds of Bayesian DP-optimal decisions and value functions, which provide practical yet computationally feasible alternatives.
Another policy that Bayesian DP literature adopts is a myopic policy, where the decision-maker optimizes inventory decisions one period at a time, solving a single-period problem without considering how the chosen order quantity impacts future learning of demand parameters. This myopic approach has been widely studied in inventory management (see (Kamath and Pakkala, 2002), (DeHoratius et al., 2008), (Bisi et al., 2011), (Besbes et al., 2022), (Chuang and Kim, 2023)). While myopic policies offer computational advantages, they often lead to suboptimal long-term inventory strategies, as they fail to fully account for the value of exploration in learning-based settings.
Specifically, we would like to compare our work with Theorem 3 in (Besbes et al., 2022). Our approach differs by benchmarking against the ground truth policy, whereas Bayesian DP-based approaches in prior work compare against Bayesian DP-optimal policies. In the frequentist setting, the ground truth policy corresponds to the true demand parameter $\theta^{*}$. In contrast, in the Bayesian setting, the policy evolves dynamically, selecting the optimal decision distribution in each round rather than following the dynamic programming approach, which sums the policy over $T$ rounds and minimizes it (as in traditional backward induction approaches). This distinction in benchmarking leads to a fundamentally different regret characterization. Unlike Bayesian DP policies, which rely on backward induction to compute the best policy in expectation, our method ensures that the regret bound scales as $\sqrt{T}$. This result highlights how our approach inherently differs in how the policies are constructed, updated, and evaluated over time. Moreover, since our benchmark does not rely on the dynamic programming framework, it avoids the computational overhead associated with backward induction, making it more scalable and efficient.
Compared to offline Bayesian DP methods, our work employs TS to learn the unknown demand parameter, providing a simpler and more computationally efficient alternative. Instead of requiring a full-state space formulation and solving for an optimal policy via backward induction, our approach dynamically learns the demand distribution while simultaneously making inventory decisions. TS offers a practical solution for real-time decision-making, as it balances exploration and exploitation without requiring predefined state transitions or explicit value function approximations.
1.3.2. Non-Parametric and Other Related Newsvendor Literature
Next, we discuss another line of research that focuses on nonparametric methods for solving joint demand estimation and inventory optimization problems. Unlike the Bayesian approach, which relies on a specific parametric demand distribution, this approach does not impose any predefined distributional assumptions on demand. Instead, it estimates demand directly from observed data. Researchers in this area develop models and algorithms that adjust inventory decisions based on demand observations without assuming a fixed functional form. For instance, (Huh and Rusmevichientong, 2009) proposes non-parametric adaptive policies that generate ordering decisions over time, allowing for flexibility in adapting to various demand patterns. Similarly, (Agrawal and Jia, 2019) proposes an updating confidence interval method that employs a phase-based UCB approach for learning and decision-making, which iteratively refines order quantities as more data becomes available. In our experiments detailed in Section 4, we demonstrate that TS outperforms these algorithms
Additionally, recent studies have explored the integration of feature-based learning into inventory systems with censored demand, introducing approaches that leverage contextual information to improve decision-making. For instance, (Ding et al., 2024) proposes the feature-based adaptive inventory algorithm and the dynamic shrinkage algorithm, which utilize observed demand patterns and additional features to dynamically adjust inventory policies. These algorithms aim to enhance the responsiveness of inventory systems to changing demand conditions by incorporating relevant external information. Meanwhile, (Tang et al., 2025) extends this idea to a pricing problem under censored demand, demonstrating how contextual features can inform pricing strategies in uncertain demand environments, thereby improving revenue management.
1.3.3. Thompson Sampling regret analysis
In this section, we highlight how our TS regret analysis differs from previous approaches, such as those in (Russo and Van Roy, 2014) and (Russo and Van Roy, 2016). Specifically, we leverage the problem structure to reformulate regret analysis in terms of the convergence of the posterior parameter, providing a more structured and interpretable framework for regret analysis. This perspective allows for a clearer understanding of how the learning process influences decision-making over time and offers insights into the dynamics of regret reduction.
A key distinction between our work and (Russo and Van Roy, 2014) lies in how exploration and exploitation are handled. Unlike UCB-based methods, which construct deterministic confidence intervals to manage the exploration-exploitation trade-off, TS operates in a Bayesian framework, dynamically updating the posterior distribution based on observed data. This posterior-driven approach allows for more adaptive decision-making, where uncertainty is reduced naturally over time without the need for explicit confidence interval constructions. By sampling from the posterior distribution, TS inherently balances the need to explore suboptimal actions to gather information and the desire to exploit actions that currently appear optimal, leading to more efficient learning and improved performance in practice.
Additionally, our analysis differs from the information-theoretic regret framework of (Russo and Van Roy, 2016), which relies on the concept of the information ratio to bound regret. While this approach has been successfully applied to fully observed bandit problems, it is not directly applicable to our setting, where demand is censored. In censored demand environments, the information ratio is difficult to compute due to missing observations on lost-sales, making the standard information-theoretic regret bounds less effective. Instead, our analysis is tailored to the specific structural properties of the newsvendor problem with censored demand, ensuring that regret is properly quantified under partial observation constraints. By focusing on the convergence properties of the posterior distribution, we provide a regret analysis that is both practical and theoretically sound in the context of censored data.
Unlike existing methods that focus on confidence-based or information-theoretic approaches, we introduce a novel regret analysis that directly links regret minimization to the convergence of the posterior distribution. This formulation offers new insights into how uncertainty reduction in the posterior translates to improved decision-making, setting the foundation for future Bayesian regret analysis in inventory and learning-based optimization problems. By establishing a direct connection between the learning dynamics of the posterior distribution and the resulting regret, our analysis provides a deeper understanding of the mechanisms driving performance in Bayesian adaptive algorithms and opens avenues for further research in this area.
The rest of the paper is organized as follows: In Section 2, we present the preliminaries and the newsvendor setup, establishing the foundation for our study. Section 2.3 details the dynamics of the TS algorithm as applied to the newsvendor problem, explaining its operation and relevance to inventory decision-making under uncertainty. In Section 3, we provide a regret analysis along with a sketch of the proof, quantifying the performance of our approach compared to the optimal benchmark. Section 4 showcases numerical experiments where we evaluate our algorithm against existing methods, highlighting its practical effectiveness. In Section 5, we discuss the broader applicability of our framework, outlining how TS with censored feedback can be implemented in other contexts. Finally, Section 6 concludes our work, summarizing findings and suggesting potential future research directions. All proofs supporting our theoretical claims are provided in the Appendix.
2. Model Setup and the Thompson Sampling Algorithm
@@
2.1. Repeated Newsvendor Model
Following the setup by (Bisi et al., 2011) and (Chuang and Kim, 2023), we consider a Repeated Newsvendor Model in which a retailer sells a single perishable product over a discrete and finite decision horizon. A Bayesian Repeated Newsvendor Model can be defined as a tuple $(T,f_{\theta_{\star}}(\cdot),\rho_{0}(\cdot),h,p)$, where $T\in\mathbb{R}^{+}$ is the known length of decision horizon, $f_{\theta_{\star}}(\cdot)$ is the known class of demand distributions, parameterized by an unknown parameter $\theta_{\star}$. We define the expression of $f_{\theta_{\star}}(\cdot)$ and $\rho_{0}(\cdot)$ in the next subsection. $h>0$ is the unit overage cost, and $p>0$ is the unit stock-out penalty. $h$ occurs if there is any leftover. $p$ occurs if there is any unmet demand.
The dynamic is defined as follows. Before the decision-making process, the parameter $\theta_{*}$ is unknown. At time $t\in[T]$, three events happens sequentially:
  1. (1)
@@ 2.1. Repeated Newsvendor Model
  2. (2)
    The demand $D_{t}$ is i.i.d generated from demand distribution $f_{\theta_{*}}(\cdot)$.
  3. (3)
@@
    (2) $g(y_{t},D_{t})=\mathbb{E}\left[h\left(y_{t}-D_{t}\right)^{+}+p\left(D_{t}-y_{t}\right)^{+}\right].$
The retailer knows the length of horizon $T$, the class of demand distributions $f_{\theta}(\cdot)$, the prior distribution $\rho_{0}$, $h$ and $p$, but does not know the exact value of $\theta_{*}$.
According to (Chuang and Kim, 2023). We denote $H=\left\{H_{t}\right\}$ the natural filtration generated by the right-censored sales data, i.e $H_{t}=\sigma\left\{(Y_{i},\delta_{i}):i\leq t\right\}$, where $Y_{t}=D_{t}\wedge y_{t}$ and $\delta_{t}=1\left[D_{t}<y_{t}\right]$. DM chooses an action $y_{t}$. The DM aims to minimize the total expected cost in the $T$- period online phase. We quantify the performance guarantee of the DM’s non-anticipatory policy $\pi$ by its regret. We define regret as $\operatorname{Regret(T,\pi,\theta_{\star})}$ be the regret with respect to a fixed $\theta_{\star}$.
Definition 2.1.
(3) $\operatorname{Regret(T,\pi,\theta_{\star})}=\mathbb{E}\left[\sum_{t=1}^{T}g\left(y_{t},D_{t}\right)-\sum_{t=1}^{T}g\left(y_{\star},D_{t}\right)\mid\theta_{\star}\right].$
For simplicity, throughout the paper we abbreviate $\operatorname{Regret(T,\pi,\theta_{\star})}$ as $\operatorname{Regret(T,\theta_{\star})}$.
$y_{*}=\mathop{\arg\max}_{y}\ \mathbb{E}\left[h\left(y-D_{t}\right)^{+}+p\left(D_{t}-y\right)^{+}\right]=F^{-1}_{\theta_{\star}}\left(\frac{p}{p+h}\right).$
We note by definition, the regret here is essentially a frequentist (non-Bayesian) regret. In this definition, $\theta_{\star}$ should be viewed as a fixed parameter. Even though our work focuses on the development of TS, which is a Bayesian online learning algorithm, our regret analysis holds for the more general case in which the prior demand can be drawn from arbitrary probability distributions.
2.2. Preliminaries

### v2

@@
Abstract
In this paper, we investigate the performance of Thompson Sampling (TS) for online learning with censored feedback, focusing primarily on the classic repeated newsvendor model–a foundational framework in inventory management–and demonstrating how our techniques can be naturally extended to a broader class of problems. We first model demand using a Weibull distribution and initialize TS with a Gamma prior to dynamically adjust order quantities. Our analysis establishes optimal (up to logarithmic factors) frequentist regret bounds for TS without imposing restrictive prior assumptions. More importantly, it yields novel and highly interpretable insights on how TS addresses the exploration-exploitation trade-off in the repeated newsvendor setting. Specifically, our results show that when past order quantities are sufficiently large to overcome censoring, TS accurately estimates the unknown demand parameters, leading to near-optimal ordering decisions. Conversely, when past orders are relatively small, TS automatically increases future order quantities to gather additional demand information. Then, we extend our analysis to general parametric distribution family and provide proof for Bayesian regret. Extensive numerical simulations further demonstrate that TS outperforms more conservative and widely-used approaches such as online convex optimization, upper confidence bounds, and myopic Bayesian dynamic programming.
1 Introduction
@@
More broadly, the repeated newsvendor problem serves as a key representative of a broader class of problems referred to as “online learning with censored feedback.” In this setting, the observation is always the minimum of an unknown random variable and the chosen action. For instance, in the newsvendor problem, the censored feedback corresponds to the minimum of the demand and the stock order. Similarly, in an auction, it is given by the minimum of the buyer’s willingness to pay and the seller’s set price. These problems inherently exhibit a trade-off between “large exploration”—choosing a sufficiently large action to better observe demand or willingness to pay for more accurate estimation—and “optimal exploitation”—making the right decision to minimize regret. While this paper primarily focuses on the repeated newsvendor problem, we also take an initial step toward systematically exploring the broader class of online learning problems with censored feedback. Existing studies on the repeated newsvendor problem have established a $\sqrt{T}$-regret bound under fairly general unknown demand distributions and censored feedback, often leveraging the online convex optimization (OCO) framework Huh and Rusmevichientong (2009). However, as widely recognized in the bandit and online learning literature Chapelle and Li (2011); Seldin and Slivkins (2014); Xu and Zeevi (2023), TS often outperforms OCO-based approaches (which were originally developed for adversarial online learning) as well as other methods such as the Upper Confidence Bound (UCB). These advantages are supported by extensive numerical experiments and theoretical analyses in the aforementioned studies, as well as in our own work. This motivates us to adopt TS as the preferred approach for the repeated newsvendor problem and beyond.
1.1 Main Contributions
@@
In this paper, we investigate an online learning problem with censored feedback using the classic newsvendor model–one of the most fundamental frameworks in inventory management–as our pivotal example. Specifically, we consider a setting where the true demand $D_{t}$ is unknown, the action $y_{t}$ is the order quantity, and the observation $Y_{t}$ is censored feedback given by $Y_{t}=\min\{D_{t},y_{t}\}.$ Where demand is exactly observed when sales are less than the order quantity, that is, when $D_{t}<y_{t}$; and the demand is censored at the order quantity when sales equal $y_{t}$, that is, when $D_{t}\geq y_{t}$. The newsvendor setting experiences censored feedback–the decision-maker never observes lost-sales if demand exceeds the order quantity. This makes it difficult to accurately estimate demand, as it requires finding the right balance between not ordering too much to prevent excess inventory and placing larger orders to better understand how much demand is actually being missed (i.e., the afore-mentioned exploration-exploitation trade-off). To address this trade-off, One the one side, Our analysis focuses on Bayesian regret analysis for general demand distribution. On the other side, We also establish frequentist guarantees under Weibull distribution–a flexible and widely used parametric family. Our key insights include:
1.1.1 Frequentist under Weibull Distribution
@@
Regardless of the algorithm used, we derive the confidence interval for demand estimation under censored feedback $Y_{t}$. The estimation error at round $t$ scales inversely with $\sum_{i=1}^{t-1}(1-e^{\theta_{\star}y_{i}^{k}})$, where $\theta_{\star}$ and $k$ are the scale and shape parameters of the Weibull distribution. This provides a rigorous quantification of how smaller past actions lead to larger errors and highlights the critical trade-off between large exploration and optimal exploitation.
@@
1.1.2 Bayesian Regret under General Parametric Families
Beyond the Weibull demand, we extend our analysis to general parametric family. To handle censored observations, we estimate the true demand CDF using the Kaplan–Meier (KM) estimator and construct a plug-in estimator that selects the closest parametric fit to the KM estimator. Those allows us to establish the a upper confidence interval for Bayesian estimator. Finally we bound the bayesian regret through Lipschitz property of the newsvendor cost.
1.1.3 Empirical Effectiveness
@@
We conduct extensive numerical experiments demonstrating that TS yields competitive performance in terms of cumulative regret, outperforming existing widely-used approaches such as OCO, UCB, and myopic Bayesian dynamic programming. These experiments confirm the widely recognized belief on the effectiveness of TS in online learning and bandit literature and practical applications Chapelle and Li (2011); Seldin and Slivkins (2014); Xu and Zeevi (2023).
1.2 Related Work
Bayesian Dynamic Programming Literature
The first stream of research formulates this problem using an offline dynamic programming (DP) approach, typically solved via backward induction. However, backward induction often suffers from the curse of dimensionality, making it computationally intractable for large-scale problems. Consequently, existing literature focuses on heuristic solutions as approximations to the optimal policy. Chen (2010) propose heuristics based on bounds of Bayesian DP-optimal decisions. Another policy that Bayesian DP literature adopts is a myopic policy, where the decision-maker optimizes inventory decisions one period at a time. This myopic approach has been widely studied in inventory management (see Kamath and Pakkala (2002), DeHoratius et al. (2008), Bisi et al. (2011), Besbes et al. (2022), Chuang and Kim (2023)). Our approach differs by benchmarking against the ground truth policy, whereas Bayesian DP-based approaches compare against Bayesian DP-optimal policies. In the frequentist setting, the ground truth policy corresponds to the true demand parameter ($\theta^{*}$). This distinction in benchmarking leads to a fundamentally different regret characterization. Unlike Bayesian DP policies, our method ensures that the regret bound scales as ($\sqrt{T}$). Compared to offline Bayesian DP methods, our work employs TS to learn the unknown demand parameter, providing a simpler and more computationally efficient alternative. Instead of requiring full-state space formulation and solving for an optimal policy via backward induction, our approach dynamically learns the demand distribution while simultaneously making inventory decisions. TS offers a practical solution for real-time decision-making, balancing exploration and exploitation without requiring predefined state transitions.
Thompson Sampling Regret Analysis
In this section, we highlight how our TS regret analysis differs from previous approaches, such as those in Russo and Van Roy (2014) and Russo and Van Roy (2016). Specifically, we leverage the problem structure to reformulate regret analysis in terms of the convergence of the posterior parameter, providing a more structured and interpretable framework for regret analysis. A key distinction between our work and Russo and Van Roy (2014) lies in how exploration and exploitation are handled. Unlike UCB-based methods, which construct deterministic confidence intervals to manage the exploration-exploitation trade-off, TS operates in a Bayesian framework, dynamically updating the posterior distribution based on observed data. By sampling from the posterior distribution, TS inherently balances the need to explore suboptimal actions to gather information and the desire to exploit actions that currently appear optimal. Additionally, our analysis differs from the information-theoretic regret framework of Russo and Van Roy (2016), which relies on the concept of the information ratio to bound regret. While this approach has been successfully applied to fully observed bandit problems, it is not directly applicable to our setting, where demand is censored. In censored demand environments, the information ratio is difficult to compute due to missing observations on lost-sales. Instead, our analysis uses the specific structural properties of the newsvendor problem with censored demand. Unlike existing methods that focus on confidence-based or information-theoretic approaches, we introduce a novel regret analysis that directly links regret minimization to the convergence of the posterior distribution.
2 Model Setup and the Thompson Sampling Algorithm
@@
2.1 Repeated Newsvendor Model
Following the setup by Bisi et al. (2011) and Chuang and Kim (2023), we consider a Repeated Newsvendor Model in which a retailer sells a single perishable product over a discrete and finite decision horizon. A Bayesian repeated newsvendor Model can be defined as a tuple $(T,f_{\theta_{\star}}(\cdot),\rho_{0}(\cdot),h,p)$, where $T\in\mathbb{R}^{+}$ is the known length of decision horizon, $f_{\theta_{\star}}(\cdot)$ is the known class of demand distributions, parameterized by an unknown parameter $\theta_{\star}$. We define the expression of $f_{\theta_{\star}}(\cdot)$ and $\rho_{0}(\cdot)$ in the next subsection. $h>0$ is the unit overage cost, and $p>0$ is the unit stock-out penalty. $h$ occurs if there is any leftover. $p$ occurs if there is any unmet demand. The dynamic is defined as follows. Before the decision-making process, the parameter $\theta_{\star}$ is unknown. At time $t\in[T]$, three events happen sequentially:
  1. 1.
@@ 2.1 Repeated Newsvendor Model
  2. 2.
    The demand $D_{t}$ is i.i.d generated from demand distribution $f_{\theta_{\star}}(\cdot)$.
  3. 3.
@@
    $g(y_{t},D_{t})=\mathbb{E}\left[h\left(y_{t}-D_{t}\right)^{+}+p\left(D_{t}-y_{t}\right)^{+}\right].$ (1)
The retailer knows the length of horizon $T$, the class of demand distributions $f_{\theta}(\cdot)$, the prior distribution $\rho_{0}$, $h$ and $p$, but does not know the exact value of $\theta_{\star}$.
According to Chuang and Kim (2023). We denote $H=\left\{H_{t}\right\}$ the natural filtration generated by the right-censored sales data, i.e $H_{t}=\sigma\left\{(Y_{i},\delta_{i}):i\leq t\right\}$, where $Y_{t}=D_{t}\wedge y_{t}$ and $\delta_{t}=1\left[D_{t}<y_{t}\right]$. The retailer chooses an action $y_{t}$. The retailer aims to minimize the total expected cost in the $T$- period online phase. We quantify the performance guarantee of the DM’s non-anticipatory policy $\pi$ by its regret. We define regret as $\operatorname{Regret(T,\pi,\theta_{\star})}$ be the regret with respect to a fixed $\theta_{\star}$.
Definition 1.
$\operatorname{Regret(T,\pi,\theta_{\star})}=\mathbb{E}\left[\sum_{t=1}^{T}g\left(y_{t},D_{t}\right)-\sum_{t=1}^{T}g\left(y_{\star},D_{t}\right)\mid\theta_{\star}\right].$
Definition 2.
$\operatorname{BayeRegret(T,\pi)}=\mathbb{E}\left[\sum_{t=1}^{T}g\left(y_{t},D_{t}\right)-\sum_{t=1}^{T}g\left(y_{\star},D_{t}\right)\right].$
For simplicity, throughout the paper we abbreviate $\operatorname{Regret(T,\pi,\theta_{\star})}$ as $\operatorname{Regret(T,\theta_{\star})}$. The optimal order quantity is given by $y_{\star}=\mathop{\arg\max}_{y}\ \mathbb{E}\left[h\left(y-D_{t}\right)^{+}+p\left(D_{t}-y\right)^{+}\right]=F^{-1}_{\theta_{\star}}\left(\frac{p}{p+h}\right).$ which corresponds to the critical quantile of the demand distribution when $\theta_{\star}$ is known.
2.2 Preliminaries